# Describing the Non-Universal Galaxy Merger Timescales in IllustrisTNG: Effects of Host Halo Mass, Baryons, and Sample Selection

### v1

@@ 1 Introduction
Galaxy mergers play a fundamental role in the hierarchical structure formation paradigm, shaping the evolution of galaxies over cosmic time (Toomre & Toomre, 1972; White & Rees, 1978; White & Frenk, 1991; Lacey & Cole, 1993). According to the $\Lambda$CDM model, galaxies grow through a combination of in-situ star formation and mergers. (Mo et al., 2010; Rodriguez-Gomez et al., 2016), making the study of mergers essential for understanding galaxy evolution. The timescale over which a merging satellite galaxy coalesces with its host influences various astrophysical processes, including star formation, black hole growth, and morphological transformation (Hopkins et al., 2006). Despite its importance, accurately predicting merger timescales remains a complex challenge due to the interplay of gravitational dynamics, baryonic physics, and environmental factors.
Early studies of galaxy mergers, primarily based on numerical simulations and analytical models, focused on the role of dynamical friction in driving satellite galaxies toward their host centers. Chandrasekhar (1943) demonstrated that the influence of random encounters on a object moving through a homogeneous and isotropic distribution of lighter objects can be effectively approximated by a dynamical friction force:
@@
However, in reality, the situation is more complex, and the assumptions used above may not always hold. For example, tidal stripping can gradually reduce the satellite’s mass, an effect not accounted for in the previous calculations. Additionally, factors such as the non-spherical density distribution of host halos, the presence of substructures, the evolution of the host halo itself, and baryonic effects further complicate the merger process and are not considered in these simplified models. Therefore, to better formulate merger timescales using only initial conditions for easier application, many studies have attempted to derive more generalized fitting formulas based on N-body or hydrodynamic simulations. These efforts aim to provide a more realistic description of galaxy merger timescales (Navarro et al., 1995; Colpi et al., 1999; Velazquez & White, 1999; Taffoni et al., 2003; Boylan-Kolchin et al., 2008; Jiang et al., 2008; Wetzel & White, 2010; Simha & Cole, 2017; Poulton et al., 2021). The generalized fitting formulas are typically expressed as
@@
where a common choice for the Coulomb logarithm is $\Lambda=(1+M_{\rm{host}}/M_{\rm{sat}})$. For example, using idealized N-body simulations, Boylan-Kolchin et al. (2008) found that $f(\epsilon)=e^{1.9\epsilon}$, with $\beta=1.3$ and $\gamma=1.0$. However, results from cosmological hydrodynamic simulations by Jiang et al. (2008) indicate a much weaker dependence of merger timescales on initial orbits, with $f(\epsilon)=0.94\epsilon^{0.6}+0.6$, $\beta=1.0$, and $\gamma=0.5$. Furthermore, it is also worth mentioning that Poulton et al. (2021) argued that, in cosmological N-body simulations, merger timescales primarily depend on the pericentric distance of satellites from the host center.
@@ 2.1 IllustrisTNG
Figure 3: Comparison of the merger timescales from TNG100-1 (top) and TNG300-1 (bottom) with the best-fit models. The first three columns compare each component: $Ae^{\alpha\epsilon}$, $(M_{\rm{host}}/M_{\rm{sat}})^{\beta}/\ln{(1+M_{\rm{host}}/M_{\rm{sat}})}$, and $(r_{c}/r_{\rm{host}})^{\gamma}$. In each case, the simulation data represents the merger timescales normalized by the best-fit models excluding the respective component. The last column compares the overall merger timescales between the simulation and the models.
@@ 2.2 Merger samples
To ensure that both central and satellite galaxies are well resolved, we select host halos with virial masses $M_{\rm{vir}}(z=0)>10^{10.7}h^{-1}M_{\odot}$ for TNG100-1 and $M_{\rm{vir}}(z=0)>10^{11.7}h^{-1}M_{\odot}$ for TNG300-1 at $z=0$. Identifying satellite galaxies that have merged into central galaxies is nontrivial because of the challenges in robustly defining mergers in simulations. A merger, as defined by the SUBLINK merger tree, can correspond to one of three scenarios: a real merger, the actual disruption of a satellite (He et al., 2025), or the artificial disruption of a satellite (van den Bosch & Ogiya, 2018). Distinguishing among these scenarios is essential for accurately studying galaxy merger timescales, since including disruption events can significantly shorten the measured durations. Therefore, following Boylan-Kolchin et al. (2008), we distinguish mergers from satellite disruptions using not only stellar mass or particle count, but also the final specific angular momentum, $j_{\rm{merger}}$, which, as shown by Boylan-Kolchin et al. (2008), is a better indicator than $r_{\rm{merger}}$. In Figure 1, we show the joint distribution of the infall stellar mass of satellite galaxies $M_{*}^{\rm{sat}}(z_{\rm infall})$ and the ratio $j_{\rm{merger}}/j_{\rm{infall}}$ of their specific angular momentum before the merger to that at infall. The stellar mass of galaxies is defined as the total stellar mass enclosed within twice the stellar half-mass radius. We show the results for both TNG100-1 and TNG300-1 for all systems with $z_{\rm{infall}}<4$. It is evident that $j_{\rm{merger}}/j_{\rm{infall}}$ decreases with increasing stellar mass. Moreover, at a fixed stellar mass, TNG100-1 exhibits, on average, lower $j_{\rm{merger}}/j_{\rm{infall}}$ values compared to TNG300-1. This suggests that galaxies with fewer particles are more susceptible to disruption, consistent with the results reported in van den Bosch & Ogiya (2018). To construct the merger sample, we apply both stellar mass and angular momentum cuts, as indicated by the red lines in Figure 1. Specifically, we use an infall stellar mass cut of $M_{*}^{\rm sat}(z_{\rm infall})>10^{8.3}h^{-1}M_{\odot}$ for satellite galaxies in TNG100-1 and $M_{*}^{\rm sat}(z_{\rm infall})>10^{9.3}h^{-1}M_{\odot}$ for those in TNG300-1, along with an angular momentum cut of $j_{\rm{merger}}/j_{\rm{infall}}<0.1$. After applying the selection criteria, we identify a total of 3,792 mergers in TNG100-1 and 22,635 mergers in TNG300-1.
@@ 2.3 Merger timescales and orbit parameters
The goal of this paper is to investigate the prediction of merger timescales for satellite galaxies based on their initial conditions at the time of infall. To achieve this, we calculate the infall and merger times, along with the orbital parameters and mass ratios at the infall time, for these satellite galaxies.
@@
Since the simulations are stored in discrete snapshots, determining the exact infall and merger times requires approximations. For the infall time, we use the positions, velocities, and times of satellite galaxies from the snapshots immediately before and after infall. By assuming a constant acceleration in the radial direction, we solve for the precise time at which the satellite galaxies cross $r_{\rm{vir}}$. For the merger time, following Jiang et al. (2008), we adopt the midpoint between the snapshots immediately before and after the merger as the merger time. While this approach may introduce some uncertainty in the precise determination of the merger time, the dense temporal sampling provided by the large number of snapshots ensures that this uncertainty represents only a minor source of error.
@@
For the initial conditions, all quantities are calculated at the snapshot immediately preceding infall. Three sets of parameters are typically required to determine the merger timescale: the mass ratio $M_{\rm{sat}}/M_{\rm{host}}$, the orbit parameters $\Theta$, and the dynamical time $T_{\rm{dyn}}$. For the mass ratio, we define $M_{\rm{host}}$ as the virial mass $M_{\rm{vir}}$ of the host halo and $M_{\rm{sat}}$ as the bound subhalo mass of the satellite. The dynamical time $T_{\rm{dyn}}$ is defined as $T_{\rm{dyn}}=\sqrt{r_{\rm{host}}^{3}/(GM_{\rm{host}})}$, where $r_{\rm{host}}$ is the virial radius $r_{\rm{vir}}$ of the host halos. For a spherical potential, the orbit of a satellite can be fully determined by its energy and angular momentum $(E,\ L)$. In practice, two dimensionless quantities are commonly used: $\epsilon$ and $r_{c}(E)/r_{\rm{host}}$. Here, $\epsilon=j_{\rm{infall}}/j_{c}(E)$ represents the circularity of the orbit, while $r_{c}(E)$ and $j_{c}(E)$ denote the radius and specific angular momentum, respectively, of a circular orbit with the same energy $E$. As in many studies (Lacey & Cole, 1993; Jiang et al., 2008, 2015), we adopt a SIS potential for halos to calculate the initial orbit parameters, as it is more convenient than the Navarro–Frenk–White (NFW; Navarro et al., 1997) potential. The differences in orbit parameters between the SIS and NFW potentials have been explored in Jiang et al. (2015, see Figure 11). While the NFW potential may be more realistic, we find that the parameters derived from the SIS potential already provide accurate predictions for merger timescales. Given its simplicity and effectiveness, we retain the use of the SIS potential in our analysis.
Table 1: Best-fit parameters of merger timescale models. The last model (infall hydro) is recommended for general use.
infall hydro any Eq. 7, 8, 10, 11, 12 0.652 12.599 0.603 1.135 0.426 0.532 0.188 0.421 0.805 60.036 -0.232
@@
In Figure 2, we show the distributions of the initial conditions $M_{\rm{sat}}/M_{\rm{host}}$, $\epsilon$, and $r_{c}(E)/r_{\rm{host}}$ for our merger samples. Despite the differences in the selection criteria for $M_{*}^{\rm{sat}}$ and $M_{\rm{vir}}$ between TNG100-1 and TNG300-1, their merger samples exhibit very similar initial distributions. Similar to Jiang et al. (2008), most mergers in our sample have a mass ratio larger than 0.03, as smaller satellites cannot merge within a Hubble time. Additionally, we find that our merger sample exhibits smaller $\epsilon$ values compared to the complete infall sample studied in Jiang et al. (2015), while the $r_{c}(E)/r_{\rm{host}}$ distribution is quite similar. This arises because radial orbits have shorter merger timescales and are more likely to be included in our merger sample. The impact of these selection effects on our results will be examined in Section 5.2 using the whole infall sample rather than only merger sample.
@@ 3.1 Modeling the entire merger samples
As in previous studies (Jiang et al., 2008; Simha & Cole, 2017; Poulton et al., 2021), we begin by attempting to construct a universal model for the entire merger sample. We adopt the merger timescale model from Boylan-Kolchin et al. (2008), which provides a better fit to our results. The merger timescale model we adopt is
$\frac{T_{\rm{merger}}}{T_{\rm{dyn}}}=Ae^{\alpha\epsilon}\frac{(M_{\rm{host}}/M_{\rm{sat}})^{\beta}}{\ln(1+M_{\rm{host}}/M_{\rm{sat}})}\left[\frac{r_{c}(E)}{r_{\rm{host}}}\right]^{\gamma}\,,$ (5)
@@
where $N$ is the number of mergers in the sample, $p$ is the number of free parameters, $T_{\rm{sim}}$ represents the merger timescales calculated from simulations, and $T_{\rm{mod}}$ denotes the timescales predicted by the model. We obtain the following best-fit parameters for TNG100-1:
@@
In Figure 3, we compare the merger timescales from simulations and best-fit models. Each component in Equation 5 is examined separately, with the simulation data showing merger timescales normalized by the best-fit models that exclude the corresponding component. We find that the dependence of merger timescales on all components—$\epsilon$, $M_{\rm{sat}}/M_{\rm{host}}$, and $r_{c}/r_{\rm{host}}$—is well captured by Equation 5. The overall scatter in both TNG100-1 and TNG300-1 is small, indicating that our model provides accurate predictions for merger timescales. However, although our model fits the merger timescales in both TNG100-1 and TNG300-1 samples well, the best-fit parameters differ, particularly for $\alpha$, which describes the dependence on orbit circularity $\epsilon$. This suggests that galaxy merger timescales may not be universal and could depend on sample properties.

### v4

# Describing the Nonuniversal Galaxy Merger Timescales in IllustrisTNG: Effects of Host Halo Mass, Baryons, and Sample Selection

@@ 1 Introduction
Galaxy mergers play a fundamental role in the hierarchical structure formation paradigm, shaping the evolution of galaxies over cosmic time (Toomre & Toomre, 1972; White & Rees, 1978; White & Frenk, 1991; Lacey & Cole, 1993). According to the $\Lambda$CDM model, galaxies grow through a combination of in-situ star formation and mergers. (Mo et al., 2010; Rodriguez-Gomez et al., 2016), making the study of mergers essential for understanding galaxy evolution. The timescale over which a merging satellite galaxy coalesces with its host influences various astrophysical processes, including star formation, black hole growth, and morphological transformation (Hopkins et al., 2006). Therefore, merger timescales are essential for semi-analytical models of galaxy formation, as it controls the merger of satellite thus determine the satellite aboundance and also the processes mentioned above (White & Frenk, 1991; Kauffmann et al., 1993; Cole et al., 2000; De Lucia & Blaizot, 2007; Guo et al., 2011; Henriques et al., 2015). However, despite its importance, accurately predicting merger timescales remains a complex challenge due to the interplay of gravitational dynamics, baryonic physics, and environmental factors.
Early studies of galaxy merger timescales, primarily based on numerical simulations and analytical models, focused on the role of dynamical friction in driving satellite galaxies toward their host centers. Chandrasekhar (1943) demonstrated that the influence of random encounters on a object moving through a homogeneous and isotropic distribution of lighter objects can be effectively approximated by a dynamical friction force:
@@
However, in reality, the situation is more complex, and the assumptions used above may not always hold. For example, tidal stripping can gradually reduce the satellite’s mass (Taylor & Babul, 2001; Delfino et al., 2022; Du et al., 2024), an effect not accounted for in the previous calculations. Additionally, factors such as the non-spherical density distribution of host halos, the presence of substructures, the evolution of the host halo itself, and baryonic effects further complicate the merger process and are not considered in these simplified models. Therefore, to better formulate merger timescales using only initial conditions for easier application, many studies have attempted to derive more generalized fitting formulas based on N-body or hydrodynamic simulations. These efforts aim to provide a more realistic description of galaxy merger timescales (Navarro et al., 1995; Colpi et al., 1999; Velazquez & White, 1999; Taffoni et al., 2003; Boylan-Kolchin et al., 2008; Jiang et al., 2008; Wetzel & White, 2010; Simha & Cole, 2017; Poulton et al., 2021). The generalized fitting formulas are typically expressed as
@@
where a common choice for the Coulomb logarithm is $\Lambda=(1+M_{\rm{host}}/M_{\rm{sat}})$. For example, using idealized N-body simulations, Boylan-Kolchin et al. (2008) found that $f(\epsilon)=e^{1.9\epsilon}$, with $\beta=1.3$ and $\gamma=1.0$. However, results from cosmological hydrodynamic simulations by Jiang et al. (2008) indicate a much weaker dependence of merger timescales on initial orbits, with $f(\epsilon)=0.94\epsilon^{0.6}+0.6$, $\beta=1.0$, and $\gamma=0.5$. Furthermore, it is worth noting that Poulton et al. (2021) argued that, in cosmological N-body simulations, merger timescales primarily depend on the pericentric distance of satellites from the host center, suggesting that tidal disruption or heating near pericenter may play a significant role in determining merger timescales (Taylor & Babul, 2001; van den Bosch et al., 2018).
@@ 2.2 Merger samples
To ensure that both central and satellite galaxies are well resolved, we select host halos with virial masses $M_{\rm{vir}}(z=0)>10^{10.7}h^{-1}M_{\odot}$ for TNG100-1 and $M_{\rm{vir}}(z=0)>10^{11.7}h^{-1}M_{\odot}$ for TNG300-1 at $z=0$, corresponding to more than 10,000 DM particles. Identifying satellite galaxies that have merged into central galaxies is nontrivial because of the challenges in robustly defining mergers in simulations. A merger, as defined by the SUBLINK merger tree, can correspond to one of three scenarios: a real merger, the actual disruption of a satellite (He et al., 2025), or the artificial disruption of a satellite (van den Bosch & Ogiya, 2018). A real merger is typically defined as the point at which a large fraction of a satellite’s particles become mixed and indistinguishable from those of the central galaxy in phase space. In contrast, actual disruption refers to the physical disintegration of a satellite before its material merges with the central galaxy. The observed stellar streams in the Milky Way are remnants of such disruptions (Malhan et al., 2018). Artificial disruption, on the other hand, occurs when a satellite is entirely disrupted due to limited mass and force resolution; such satellites would remain intact in higher-resolution simulations. Distinguishing among these scenarios is essential for accurately studying galaxy merger timescales, since including artificial disruption events can significantly shorten the measured durations. Since distinguishing between real and artificial disruption is challenging, we first focus on genuine merger events. Results that include disruption cases will be discussed in Section 5.2. Therefore, following Boylan-Kolchin et al. (2008), we distinguish mergers from satellite disruptions using not only stellar mass or particle count, but also the final specific angular momentum, $j_{\rm{merger}}$, which, as shown by Boylan-Kolchin et al. (2008), is a better indicator than $r_{\rm{merger}}$. In Figure 1, we show the joint distribution of the infall stellar mass of satellite galaxies $M_{*}^{\rm{sat}}(z_{\rm infall})$ and the ratio $j_{\rm{merger}}/j_{\rm{infall}}$ of their specific angular momentum before the merger to that at infall. The stellar mass of galaxies is defined as the total stellar mass enclosed within twice the stellar half-mass radius. We show the results for both TNG100-1 and TNG300-1 for all systems with $z_{\rm{infall}}<4$. It is evident that $j_{\rm{merger}}/j_{\rm{infall}}$ decreases with increasing stellar mass. Moreover, at a fixed stellar mass, TNG100-1 exhibits, on average, lower $j_{\rm{merger}}/j_{\rm{infall}}$ values compared to TNG300-1. This suggests that galaxies with fewer particles are more susceptible to disruption, consistent with the results reported in van den Bosch & Ogiya (2018). To construct the merger sample, we apply both stellar mass and angular momentum cuts, as indicated by the red lines in Figure 1. Specifically, we use an infall stellar mass cut of $M_{*}^{\rm sat}(z_{\rm infall})>10^{8.3}h^{-1}M_{\odot}$ for satellite galaxies in TNG100-1 and $M_{*}^{\rm sat}(z_{\rm infall})>10^{9.3}h^{-1}M_{\odot}$ for those in TNG300-1, along with an angular momentum cut of $j_{\rm{merger}}/j_{\rm{infall}}<0.1$. This stellar mass cut ensures that infalling subhalos contain more than 200 star particles and 5,000 dark matter particles, guaranteeing that the subhalos are well resolved and that baryonic effects are present. After applying the selection criteria, we identify a total of 3,792 mergers in TNG100-1 and 22,635 mergers in TNG300-1.
@@ 2.3 Merger timescales and orbit parameters
Merger timescales refer to the time between a satellite’s infall into the host halo and its eventual disruption or merger with the central galaxy. The goal of this paper is to investigate the prediction of merger timescales for satellite galaxies based on their initial conditions at the time of infall. To achieve this, we calculate the infall and merger times, along with the orbital parameters and mass ratios at the infall time, for these satellite galaxies.
@@
Since the simulations are stored in discrete snapshots, determining the exact infall and merger times requires approximations. For the infall time, we use the positions, velocities, and times of satellite galaxies from the snapshots immediately before and after infall. Assuming constant acceleration in the radial direction, we solve for the precise time at which the satellite galaxies first cross $r_{\rm{vir}}$ of their host halos. For the merger time, following Jiang et al. (2008), we adopt the midpoint between the snapshots immediately before and after the merger as the merger time. The dense sampling provided by the large number of snapshots results in a time interval of $0.1-0.2\,\rm{Gyr}$ between snapshots, which only significantly affects the rare, very fast merger events.
Figure 3: Comparison of the merger timescales from TNG100-1 (top) and TNG300-1 (bottom) with the best-fit models. The first three columns compare each component: $Ae^{\alpha\epsilon}$, $(M_{\rm{host}}/M_{\rm{sat}})^{\beta}/\ln{(1+M_{\rm{host}}/M_{\rm{sat}})}$, and $(r_{c}/r_{\rm{host}})^{\gamma}$. In each case, the simulation data represents the merger timescales normalized by the best-fit models excluding the respective component. The last column compares the overall merger timescales between the simulation and the models. The figures are color-coded by the number of merger events.
@@
For the initial conditions, all quantities are calculated at the snapshot immediately preceding infall. Three sets of parameters are typically required to determine the merger timescale: the mass ratio $M_{\rm{sat}}/M_{\rm{host}}$, the orbit parameters $\Theta$, and the dynamical time $T_{\rm{dyn}}$. For the mass ratio, we define $M_{\rm{host}}$ as the virial mass $M_{\rm{vir}}$ of the host halo and $M_{\rm{sat}}$ as the bound subhalo mass of the satellite. The dynamical time $T_{\rm{dyn}}$ is defined as $T_{\rm{dyn}}=\sqrt{r_{\rm{host}}^{3}/(GM_{\rm{host}})}$, where $r_{\rm{host}}$ is the virial radius $r_{\rm{vir}}$ of the host halos. For a spherical potential, the orbit of a satellite can be fully determined by its energy and angular momentum $(E,\ L)$. In practice, two dimensionless quantities are commonly used: $\epsilon$ and $r_{c}(E)/r_{\rm{host}}$. Here, $\epsilon=j_{\rm{infall}}/j_{c}(E)$ represents the circularity of the orbit, while $r_{c}(E)$ and $j_{c}(E)$ denote the radius and specific angular momentum, respectively, of a circular orbit with the same energy $E$. As in many studies (Lacey & Cole, 1993; Jiang et al., 2008, 2015), we adopt a SIS potential for halos to calculate the initial orbit parameters, as it is more convenient than the Navarro–Frenk–White (NFW; Navarro et al., 1997) potential. The differences in orbit parameters between the SIS and NFW potentials have been explored in Jiang et al. (2015, see Figure 11). While the NFW potential may be more realistic, we find that the parameters derived from the SIS potential already provide accurate predictions for merger timescales, as indicated by $\sigma_{\log}$ below. Given its simplicity and effectiveness, we retain the use of the SIS potential in our analysis.
@@
In Figure 2, we show the distributions of the initial conditions $M_{\rm{sat}}/M_{\rm{host}}$, $\epsilon$, and $r_{c}(E)/r_{\rm{host}}$ for our merger samples. Despite the differences in the selection criteria for $M_{*}^{\rm{sat}}$ and $M_{\rm{vir}}$ between TNG100-1 and TNG300-1, their merger samples exhibit very similar initial distributions. Similar to Jiang et al. (2008), most mergers in our sample have a mass ratio larger than 0.03, as smaller satellites cannot merge within a Hubble time. Additionally, we find that our merger sample exhibits smaller $\epsilon$ values compared to the complete infall sample studied in Jiang et al. (2015), while the $r_{c}(E)/r_{\rm{host}}$ distribution is quite similar. This arises because radial orbits have shorter merger timescales and are more likely to be included in our merger sample. The impact of these selection effects on our results will be examined in Section 5.2 using the full infall sample rather than only merger sample.
@@ 3.1 Modeling the entire merger samples
As in previous studies (Jiang et al., 2008; Simha & Cole, 2017; Poulton et al., 2021), we begin by attempting to construct a universal model for the entire merger sample. We adopt the exponential form of $f(\epsilon)$ from Boylan-Kolchin et al. (2008), as it provides a better fit to our results than alternative forms such as power laws. The merger timescale model we adopt is
$\frac{T_{\rm{merger}}}{T_{\rm{dyn}}}=Ae^{\alpha\epsilon}\frac{(M_{\rm{host}}/M_{\rm{sat}})^{\beta}}{\ln(1+M_{\rm{host}}/M_{\rm{sat}})}\left[\frac{r_{c}(E)}{r_{\rm{host}}}\right]^{\gamma}\,.$ (5)
Here, $\alpha$, $\beta$, and $\gamma$ control the dependence of merger timescales on orbital circularity, mass ratio, and orbital energy, respectively, while $A$ serves as a normalization factor.
@@
where $N$ is the number of mergers in the sample, $p$ is the number of free parameters, $T_{\rm{sim}}$ represents the merger timescales calculated from simulations, and $T_{\rm{mod}}$ denotes the timescales predicted by the model. The $\chi^{2}$ is minimized using the Levenberg–Marquardt algorithm. We obtain the following best-fit parameters for TNG100-1:
@@
In Figure 3, we compare the merger timescales from simulations and best-fit models. Each component in Equation 5 is examined separately, with the simulation data showing merger timescales normalized by the best-fit models that exclude the corresponding component. We find that the dependence of merger timescales on all components—$\epsilon$, $M_{\rm{sat}}/M_{\rm{host}}$, and $r_{c}/r_{\rm{host}}$—is well captured by Equation 5. The overall scatter in both TNG100-1 and TNG300-1 is small with $\sigma_{\log}$ smaller than 0.15, indicating that our model provides accurate predictions for merger timescales. However, although our model fits the merger timescales in both TNG100-1 and TNG300-1 samples well, the best-fit parameters differ, particularly for $\alpha$, which describes the dependence on orbit circularity $\epsilon$. This suggests that galaxy merger timescales may not be universal and could depend on sample properties.
Table 1: Descriptions and best-fit parameters of the four merger timescale models constructed in this study, based on different samples, starting radii, and model parameterizations. The first two rows use the merger sample starting at $r_{\rm vir}$, from baryonic and DMO runs respectively. The third uses the merger sample with arbitrary starting radii in baryonic runs. The fourth, recommended for general use, is based on the full infall sample with arbitrary starting radii in baryonic runs.
infall hydro any Eq. 7, 8, 10, 11, 12 0.323 12.484 0.555 0.657 0.328 0.497 0.326 0.408 1.077 30.890 -0.026